OpenStudy (hellokitty17):

Choose the correct simplification of (6x4y2z3)(3x4y3z5). a.18x16y6z15 b.18x8y5z8 c.9x8y5z8 d.9x16y6z15

OpenStudy (phi):

\[ 6\cdot x^4y^2z^3\cdot 3x^4y^3z^5 \] reorder (when you multiply you can change the order) \[ 6\cdot 3\cdot x^4x^4y^2y^3z^3z^5 \] the fast way to multiply is *add the exponents* if the terms have the *same base* the long way is to expand: for example x^4 means x*x*x*x and x^4 * x^4 means x*x*x*x * x*x*x*x now count the number of x's and write it in "short hand" x with a little number showing the number of x's